Please help me prove:
-a = (-1)*a

Thank you.

Answer by stanbon(75887)

-a = (-1)*a
---
You need to show the (-1)a is the additive inverse of "a".
-------------------------
a + (-1)a = (1+(-1))a ; Distributive Law
= 0*a = 0 ; Definition of additive inverse.
---
Therefore (-1)a is the additive inverse of a ; Definition of additive inverse.
-------------------------
Therefore (-1)a = -a ; Def. of equality
==========================================
Cheers,
Stan H.
==================